OpenStudy (anonymous):

secθ secθ/ tanθ+cotθ = sin^2θ

OpenStudy (jamesj):

This identity looks wrong. When θ = pi/4, sec^2 θ = 2 and hence secθ secθ/ (tanθ+cotθ) = 2/(1+1) = 1 But sin^2θ = 1/2

OpenStudy (abb0t):

NOTE: \[\sec(\theta) = \frac{ 1 }{ \cos(\theta) }, \tan(\theta)=\frac{ \sin(\theta) }{ \cos(\theta) }, \cot(\theta)=\frac{ \cos(\theta) }{ \sin(\theta) }\]